### v2

# atom sep=2em

Abstract
Analyzing synthesis pathways for target molecules in a chemical reaction network annotated with information on the kinetics of individual reactions is an area of active study. This work presents a computational methodology for searching for pathways in reaction networks which is based on integer linear programming and the modeling of reaction networks by directed hypergraphs. Often multiple pathways fit the given search criteria. To rank them, we develop an objective function based on physical arguments maximizing the probability of the pathway. We furthermore develop an automated pipeline to estimate the energy barriers of individual reactions in reaction networks. Combined, the methodology facilitates flexible and kinetically informed pathway investigations on large reaction networks by computational means, even for networks coming without kinetic annotation, such as those created via generative approaches for expanding molecular spaces. To demonstrate the methodology, we apply it on a chemical reaction network modeling the molecular space of glycolonitrile chemistry, where we search for pathways to glycine and glycolic acid using small molecules as precursors.
I Introduction
Kinetic studies of individual reactions have provided valuable insight into their mechanisms, which in turn can generate ways to optimize the reaction to suit the overall synthesis goals. However, molecules are often entangled in various concurrent reactions, which together form an interconnected web of reactions called a reaction network. Theoretical and computational modeling offers a possibility for analyzing these systems to various degrees of accuracy and computational efficiency, depending on the level of abstraction employed. In this work, we propose a method for integrating kinetics into the search for pathways in reaction networks modeled on a level of abstraction where the network is represented by a hypergraph and individual reactions are represented by directed hyperedges. The nodes of the hypergraph are molecular graphs (with nodes representing atoms and edges representing bonds) and hence the modeling is atomistically explicit. The work is a continuation of our previous work [1], which focused on the thermodynamics of individual reactions in the network. However, despite the name, thermodynamics does not help to elucidate the ‘dynamics’ of the system: how fast does the pathway in question run and how quickly does the system equilibrate? For example, under atmospheric conditions, the conversion of diamond to graphite is thermodynamically favorable. However, the energy barrier for this conversion is high, so, thankfully, for all practical purposes ‘diamonds are forever’. Therefore, kinetic parameters of reactions, such as the free energy barrier or the rate constant, might provide better alternatives for characterizing pathways.
However, before one can use kinetic parameters to search for pathways in a reaction network, the network itself has to be constructed. This can be done in various ways. In one approach, the potential energy surface of a set of molecules is explored to elucidate trajectories with low-lying saddle points, separating the reactants from the products. Computationally intensive quantum mechanical calculations are used in most cases [2, 3, 4], along with faster but less accurate semi-empirical methods in some [5, 6]. Recently, generative diffusion models have also been used to perturb molecules to generate reaction networks with unknown mechanisms [7, 8]. A different approach is to use reaction templates for iteratively expanding the network from a set of starting molecules, possibly in conjunction with filters to exclude specific reactions [9, 10, 11]. Further approaches include the generation of the reaction network of interest in a one-shot heuristical process by an expert [12, 13, 14], or the loading of an existing network from a database [15, 16].
Except for the approach where the reaction network is generated by the exploration of the potential energy surface, individual reactions will need to be annotated with the kinetic parameters from a separate source, such as databases [17], estimations informed by experiments [18, 19], or dedicated quantum mechanical calculations [20]. Empirical relations, deduced via machine learning on available data on kinetics of reactions, have been used to estimate kinetic parameters for reactions [21, 22, 23, 24, 25, 26, 27]. However, these methods have limited success on ‘out of class’ molecules [28].
Once a reaction network has been constructed and the reactions annotated with appropriate kinetic data, one can employ diverse strategies to search and analyze pathways in it [29, 30, 31]. Previous work include Beam search based on suitable crafted scoring functions utilizing priority queues [32], as well as Monte-Carlo search trees extended with a method to sample efficient pathways [33]. Some analytical strategies have been used in the study of dynamics of pathways in reaction networks [34, 35], but they are often limited by the availability of kinetic data and the scale of the networks.
In this work, we develop an objective measure to rank pathways in a reaction network maximizing the probability of the entire pathway based on physical principles. Additionally, based on the notion that energy barriers of reactions are the determining factor for the kinetics of the network, we develop an automated methodology for computing such energy barriers, building on a combination of an array of existing computational tools. A key novelty is that in contrast to the standard manual intervention required to generate reaction trajectories from which the reaction barrier can be estimated, we try to automate this part of the process. Lastly, we formulate an integer linear program which allows the user to specify queries for pathways in the network in a very flexible and generic way. Solving this integer linear program returns all pathways fulfilling the criteria, ranked by the objective measure mentioned above. In short, this gives an automated, general, and flexible methodology for analyzing reaction networks, a methodology which offers a speed-up over manual search for pathways but with comparable quality. To exemplify its utility, we apply it to a reaction network expanded from the molecules glycolonitrile, ammonia and water, where we query for pathways to glycine, an amino acid, and glycolic acid, a molecule encountered in photorespiration in plants.
II Definitions
In this section, we present the mathematical frameworks used to model reaction networks and to specify searches for pathways in them, namely directed hypergraphs and integer hyperflows. This modeling is from [36] and [37].
II.1 Directed Hypergraphs
A reaction network consists of a set of molecules, $V$, and a set of reactions, $E$. Each reaction $e\in E$ generally involves multiple molecules and can be expressed as follow [38]
$\sum_{v\in V}s_{ve}^{-}v\longrightarrow\sum_{v\in V}s_{ve}^{+}v,$
where $s_{ve}^{-}$ and $s_{ve}^{+}$ are stoichiometric coefficients indicating the number of molecules of $v$ consumed and produced, respectively.
A reaction network can be effectively modeled as a directed multi-hypergraph [37], denoted as $\mathcal{H}=(V,E)$, where vertices $v\in V$ represent molecules and hyperedges $e\in E$ represent reactions. Each hyperedge $e\equiv(e^{-},e^{+})$ consists of two bags (multi-sets) of molecules: $e^{-}$ for reactant molecules and $e^{+}$ for product molecules. The count of a molecule $v$ in the reactant bag and in the product bag for a reaction $e$ is given by $s^{-}_{ve}$ and $s^{+}_{ve}$, respectively.  [^1] This notation differs from [37] by swapping + and -.
Figure 1 illustrates a hypergraph with eight vertices and nine hyperedges. It has the one-to-one reaction $e_{0}\equiv(\{v_{0}\},\{v_{1}\})$, the many-to-one reactions $e_{3}\equiv(\{v_{0},v_{1}\},\{v_{4}\})$ and $e_{1}\equiv(\{v_{2},v_{2}\},\{v_{1}\})$, the one-to-many reaction $e_{2}\equiv(\{v_{0}\},\{v_{3},v_{3}\})$, and the many-to-many reaction $e_{4}\equiv(\{v_{1},v_{2}\},\{v_{4},v_{5}\})$.
Figure 1: A directed hypergraph (top-left), where circles are vertices and the squares are hyperedges. The other three figures show possible integer hyperflows in the hypergraph, each representing a pathway, where hyperedges with non-zero flow are drawn in bold. The inflow and the outflow are depicted as arrows in and out of the source and target vertices of the pathway. Hyperedges with multiple copies of a vertex as source or target are depicted with parallel arrows. See for instance $e_{2}$, which has $v_{3}$ twice as a target and thus represents the reaction $v_{0}\rightarrow 2v_{3}$.
II.2 Integer Hyperflows
A pathway is a sequence of reactions that converts source molecules into target molecules. To formalize pathways, [37] introduced the concept of an integer hyperflow $f$ in a hypergraph $\mathcal{H}=(V,E)$.
An integer hyperflow assigns a non-negative integer to each edge $e\in E$ which specifies the number of times that the reaction represented by $e$ is used in the pathway. A pathway effects an overall reaction with a set of source molecules and and a set of target molecules. To model these sets of source and target molecules, [37] defines auxiliary half-edges for each $v\in V$: $e_{v}^{+}=(\emptyset,v)$ specifies the role of $v$ as a source molecule and $e_{v}^{-}=(v,\emptyset)$ specifies the role of $v$ as a target molecule. Setting
$\begin{array}[]{lcl}E^{+}&=&\{e_{v}^{+}\mid v\in V\},\\ E^{-}&=&\{e_{v}^{-}\mid v\in V\},\\ \overline{E}&=&E\cup E^{+}\cup E^{-},\\ \end{array}$
the hypergraph is extended to ¯$\overline{\mathcal{H}}=(V,\overline{E})$ and an integer hyperflow is defined as a vector $f\in\mathbb{N}_{0}^{|\overline{E}|}$ satisfying the flow conservation constraint:
$\forall v\in V\colon\sum_{e\in\overline{E}}s_{ve}^{+}f_{e}-\sum_{e\in\overline{E}}s_{ve}^{-}f_{e}=0,$ (1)
where $s_{ve}^{-}$ and $s_{ve}^{+}$ are the stoichiometric coefficients from Section II.1.
Figure 1 illustrates three example hyperflows in a hypergraph, with unboxed numbers showing flows on half-edges and boxed numbers indicating flows on $E$. Only half-edges with non-zero flow are displayed.
Integer hyperflows offer a versatile way to define pathways by constraining flow on half-edges. As examples, one can specify an overall reaction by fixing flows for source and target molecules, or one can set upper and lower bounds for these flows while allowing specified sets of by-products (potentially also with bounds on their outflow). Overall, this modeling allows the problem of searching for pathways in reaction networks to be transformed into the problem of finding a flow vector $f$ satisfying Equation (1) plus a number of specified constraints. Given the integer nature of hyperflows, the language of Integer Linear Programming (ILP) is a natural framework for expressing this integer hyperflow modeling, which allows the search for the specified pathways to be carried out by any of the many ILP-solvers available. This approach is part of what the software library MØD [39] implements.
III Methods
In this section, we detail our proposed methodology for automating the search for kinetically informed pathways in reaction networks. In Sections III.1 and III.2, we first propose an objective function for ranking pathways in a reaction network annotated with kinetic data in the form of energy barriers for each reaction. In Section III.3, we then formulate an integer linear program (ILP) representing such networks and allowing the specification and execution of pathway searches. This part builds on the modeling described in Section II. We also describe how to search for not only one, but multiple solutions to a given pathway query, ordered by an objective function such as the one presented in Section III.2. Lastly in Section III.4, we describe our computational workflow for estimating the energy barriers of reactions, a workflow which enables the automatic annotation of reaction networks with kinetic data.
III.1 From Rate Constants to Edge Weights
Irrespective of whether one uses Arrhenius’ theory [40, 41], collision theory [42] or transition state theory [43, 44], the dependence of the rate constant $k_{e}$ of a reaction $e$ on the free energy barrier $G_{e}$ follows that of a negative exponential. The expression for the rate constant by Eyring’s equation from transition state theory is
$k_{e}=\frac{k_{b}T}{h}\exp\left(-\frac{G_{e}}{RT}\right)$ (2)
where $R$ is the universal gas constant, $k_{b}$ is the Boltzmann constant, and $h$ is the Planck constant, each in their appropriate units. In a well-stirred system where all the reactions in the reaction network are assumed to be concurrently taking place, the temperature can be considered to be the same for all reactions. Hence, it is not a factor affecting the relative rates of the reactions in the system.
Since our modeling is not aimed at capturing temporal changes in the concentrations of molecules, we follow [31] in making the simplifying assumption that all molecules in the reaction network are present with unit concentrations. Under this assumption, the rate constant $k_{e}$ can be interpreted as being proportional to the probability of reaction $e$ happening in a fixed small volume and time interval. Also following [31], we for a given reaction network normalize these values to sum to one:
$p(e)$ $=\frac{k_{e}}{\sum_{i\in E}k_{i}}$
$=\frac{\frac{k_{b}T}{h}\exp\left(-\frac{G_{e}}{RT}\right)}{\sum_{i\in E}\frac{k_{b}T}{h}\exp\left(-\frac{G_{i}}{RT}\right)}$ (using 2)
$=\frac{\exp\left(-\frac{G_{e}}{RT}\right)}{\sum_{i\in E}\exp\left(-\frac{G_{i}}{RT}\right)}$
$=\frac{\exp\left(-\frac{G_{e}}{RT}\right)}{D}$ (3)
Here, the denominator $\sum_{i\in E}\exp(-\frac{G_{i}}{RT})$ is the same for all reactions $e$ and has been shortened to $D$. In this paper, we assign the measure $p(e)$ as the weight of the hyperedge representing $e$. However, the user is free to assign any other measures of weight to the hyperedges in the reaction network, if deemed better suited for a particular pathway search objective.
III.2 Formulating an Objective Function
We are interested in pathways from the given source molecules to the desired target molecule(s) that have a high overall probability of occurring in practice. Considering individual reactions in a pathway as independent events (and abusing the notation $e$ to mean the event of reaction $e$ happening), the probability $p_{\text{total}}$ of the entire set of reactions in the pathway occurring in one unit of time and one unit of volume is
$p_{\text{total}}$ $=p\left(\bigwedge_{e:f_{e}>0}\left(e\wedge e\wedge\ldots f_{e}\text{ times}\right)\right)$
$=\prod_{e:f_{e}>0}p(e)^{f_{e}}$
$=\prod_{e:f_{e}>0}\left[\frac{\exp\left(-\frac{G_{e}}{RT}\right)}{D}\right]^{f_{e}}$ (using 3)
$=\frac{\prod_{e:f_{e}>0}\exp\left(-\frac{f_{e}\cdot G_{e}}{RT}\right)}{\prod_{e:f_{e}>0}D^{f_{e}}}$
$=\left.\exp\left(-\frac{\sum_{e\in E}f_{e}\cdot G_{e}}{RT}\right)\middle/D^{\sum_{e\in E}f_{e}}\right.$
Since the energy barrier $G_{e}$ is weighted by the flow $f_{e}$, the conditional sum $e:f_{e}>0$ can be substituted by a simple sum over all hyperedges $e\in E$ because hyperedges with $f_{e}=0$ do not contribute to the value of the sum. The expression for the total probability of a pathway is a fraction and we take the logarithm of it to simplify it and make it a linear expression.
$\log p_{\text{total}}$ $=\log\left[\exp\left(-\frac{\sum_{e\in E}f_{e}\cdot G_{e}}{RT}\right)\middle/D^{\sum_{e\in E}f_{e}}\right]$
$=-\frac{\sum_{e\in E}f_{e}\cdot G_{e}}{RT}-\left(\sum_{e\in E}f_{e}\right)\cdot\log D$
The last line is proportional to the following expression obtained by multiplying by $RT$
$-\sum_{e\in E}f_{e}\cdot G_{e}-RT\cdot\log D\cdot\sum_{e\in E}f_{e}$
The probability of the set of reactions in the pathway is then maximized by maximizing the value of the obtained expression:
$\max\left(-\sum_{e\in E}f_{e}\cdot\left(G_{e}+RT\cdot\log D\right)\right)$
This is equivalent to the following minimization problem:
$\min\left(\sum_{e\in E}f_{e}\cdot\left(G_{e}+RT\cdot\log D\right)\right)$ (4)
This proposed objective value (4) is a linear expression because the $G_{e}$’s are constants from the input instance, $T$ is assumed constant, and $R$ is a physical constant. Also, the value $\log D=\log\left(\sum_{i\in E}\exp\left(-\frac{G_{i}}{RT}\right)\right)$ is constant when considering a fixed reaction network.
We note that via $D$, the objective value (4) of a pathway will depend on the underlying reaction network in which it is considered embedded. This may be regarded as a weakness (also shared by [31]). On the other hand, the objective value (4) is intuitively appealing by being composed of two natural components: Its first component $\sum_{e\in E}f_{e}\cdot G_{e}$ favors pathways with small sums of energy barriers of the reactions in the pathway, weighted by the number of times it is used. Its second component $\sum_{e\in E}f_{e}\cdot\left(RT\,\log D\right)$ favors pathways with fewer reactions. The factor $RT\log D$ can be viewed as weighting the relative importance of the two components.
III.3 ILP Modeling
In this section, we express the modeling of Section II as an integer linear program (ILP). Recall the key idea that in a reaction network, pathways can be expressed as hyperflows and pathway queries can be expressed as constraints on hyperflows. Such constraints can simply be added to the ILP, after which the pathway queries can be executed by running an ILP-solver on the final ILP. We also describe how to search for multiple, structurally different solutions to a given pathway query, ordered by an objective function such as (4).
Thus, in the ILP a pathway is represented by a vector of non-negative [^2] Non-negative flow values require reversible reactions to be modeled as two separate hyperedges in the reaction network, but it also allows the flexibility to model some reactions as irreversible. integer-valued variables $f_{e}$, one variable for each reaction $e$ in the reaction network. These flow variables are subject to the key constraint (1), as well as to the user-supplied constraints that specify the pathway query under consideration.
For each reaction $e$, we add a boolean indicator variable $z_{e}$ which keeps track of whether $e$ is used in the pathway or not:
$f_{e}=0$ $\iff z_{e}=0$ (5)
$f_{e}\geq 1$ $\iff z_{e}=1$ (6)
We will use these indicator variables later when listing multiple, structurally different solutions to a given pathway query. [^3] The indicator variables can also increase expressiveness when specifying pathway queries. Many ILP solvers support addition of such indicator constraints (5) and (6) directly as implications to an ILP problem (which are then solved by a nonlinear, nonconvex reformulation of the indicator constraint [45]). If this feature is not available, one can instead use the method of linearizing these implication constraints via a large constant $M$ [46].
In summary, the resulting ILP looks as follows:
$\min\left(\sum_{e\in E}\left(f_{e}\cdot G_{e}+RT\cdot\log D\cdot f_{e}\right)\right)$ (4)
$\forall e\in E\colon$ $0\leq f_{e}$ (7)
$\forall v\in V\colon$ $\sum_{e\in\overline{E}}s_{ve}^{+}f_{e}-\sum_{e\in\overline{E}}s_{ve}^{-}f_{e}=0$ (1)
Constraints specifying the pathway query
$f_{e}=0\iff z_{e}=0$ (5)
$f_{e}\geq 1\iff z_{e}=1$ (6)
An overview of the variables and constants in the ILP can be found in Table 1 and 2, respectively.
ILP variable Datatype Description
$f_{e}$ integer Flow variable for a hyperedge $e$, non-negative
$z_{e}$ boolean Indicator variable denoting if a hyperedge $e$ is used in the flow
Table 1: Variables in the proposed ILP formulation.
Constant Datatype Description
$G_{e}$ float Weight of the hyperedge $e$
$s_{ve}^{+}$ integer Count of molecule $v$ on the LHS of a reaction $e$
$s_{ve}^{-}$ integer Count of molecule $v$ on the RHS of a reaction $e$
$RT$ float Product of universal gas constant and temperature
$M$ integer Big integer used for linearizing the constraints
(type promoted to float, if required)
Table 2: Constants in the proposed ILP formulation.
Many solvers provide a solution pool to enumerate multiple solutions from. However, one often want solutions which correspond to structurally different pathways in the reaction network. We define this as each solution using a different set of reactions. To achieve this, we iteratively add constraints to the ILP model, which forces the next solution to have a set of reactions disjoint from the previous solutions. Let $S_{i-1}$ where $i\geq 2$, be the set of hyperedges in the $i-1$ iterative solution of the ILP. In the $i$ iteration, the following constraint is added to the ILP model,
$\sum_{e\in S_{i-1}}z_{e}\leq|S_{i-1}|-1$ (8)
Iteratively adding constrains like constraint (8) ensures that the $i$-th solution does not contain any of the previous $i-1$ solutions as a subset.
III.4 Computational Framework
Above in Section III.3, we have described our method for searching for pathways in a reaction network annotated with energy barriers for each reaction. In this section, we develop a computational workflow for estimating the energy barriers of reactions, a workflow aimed at automatic annotation of reaction networks with kinetic data.
III.4.1 Method for Estimating Geometry of the Molecules
We represent individual molecules in the reaction network as molecular graphs. Vertices represent atoms and are labeled by their element name, while edges represent bonds and are labeled by their bond type (single, double, triple or aromatic bond). Most tools for estimating the energy barriers of reactions require the three-dimensional embedding of the involved molecules with coordinates for each atom. We generate these molecular embeddings using OpenBabel [47] and then refine it using xTB [48] in the following manner:
  • The molecular graph is loaded into openBabel [47] as a molecule object.
  • An initial guess for the three-dimensional embedding of the molecule is generated based on a combination of elementary rules and frequently encountered fragments by using the make3D() [49] function of openBabel. However, often these guessed structures can have substructures with steric clashes or high strain in them and therefore the structures need to be optimized.
  • A systematic search for a lower energy conformer is performed by altering the torsional angles and estimating the energy of the resulting conformer using an elementary force field method (the universal force field (UFF) [50] along with the SystematicRotorSearch() [51] function of openBabel.
  • The coordinates of each atom in this resulting low energy conformer is written out in a format which can be fed as input to xTB [48].
  • Optimization of the generated geometry is performed by xTB [52] until the structure generated by its gradient descent based on the forces on individual atoms has converged.
III.4.2 Method for Estimating Energy Barriers for the Reactions
The pipeline described above generates the coordinates for the atoms in an individual molecule. However, this data cannot be directly used to determine the energy barriers for reactions involving those molecules because one needs to embed the molecules with proper orientations to form a complex (a molecular graph with multiple connected components). Furthermore, the embedding of the complexes on both sides of the reaction have to be aligned such that interpolated atomic coordinates in intermediate configurations do not pass through an artificial high-energy arrangement. We implement the following steps for calculating the energy barriers:
  • The coordinates of the molecules on each side of a reaction are read by RDKit [53] to create a collection of molecules. A complex containing these molecules is generated by aligning entire molecules so that the reactive centers are close to each other and in proper spatial orientation. The geometries of individual connected components are left unchanged.
  • The embedding is refined by xTB to take into account the inter-molecular interactions, using semi-empirical methods. (For reactions with a single molecule on one of the sides, the optimized structure from the previous section is used directly for that molecule, skipping this step and the previous step.)
  • For the energy barrier for the reaction to be estimated correctly, the order of the atoms in files containing the coordinates of the atoms for the endpoints must be the same. Our method of generating the reaction network (see Section IV) allows us to map corresponding atoms when applying a reaction template to a substructure match in a molecular graph [54]. This atom map is used to number the atoms on the two sides of the reaction correspondingly.
  • The nudged elastic band [55] method, implemented in the Atomic Simulation Environment [56], is used to determine the energy barrier for a reaction  [^4] The nudged elastic band is a double-ended method, which is suitable in our case because both sides of the reaction are known beforehand, in contrast to single-ended methods where only the reactants are known beforehand.. A number of intermediate images are generated by interpolating the coordinates of the atoms on the two sides of the reaction for an initial trajectory guess.
  • In previous work [57], a polarizable atom interaction neural network [58] was trained on a dataset of molecular configurations around reaction trajectories, Transition1x [59], to estimate the potential energy of the configuration and the forces on the atoms. This pre-trained neural net from [57] we use to estimate the forces on the atoms (given by the gradient of potential) in each intermediate image. By the nudged elastic band method, the computed forces on individual atoms are used to nudge their position to form a lower energy configuration. Iteration of this step lowers the energy of the images, leading to a reaction trajectory with lower energy barriers along its way. When the maximum of the forces on individual atoms falls below a given threshold, the procedure is stopped.
  • A cubic spline interpolation of the energies of the intermediate images is performed. The energy barrier for the reaction is read off as the difference between the maximum along this interpolation and its starting point.
We note that the user of course is free to substitute the above workflow with other procedures for estimating energy barriers for the reactions of the network, or to read values directly from some database, all depending on the desired accuracy of modeling, the available data, and the available computational resources. The crux of our proposed methodology is that it is automated and fairly fast, which makes it applicable in situations where the available data and computational resources are inadequate, such as when using generative approaches for expanding molecular spaces.
IV Results
In this section, we apply the methods from Section III to a reaction network involving glycolonitrile (MeCN), ammonia (NH3), and water (H2O), a network previously studied in [60]. The network was expanded using the graph transformation based generative framework MØD, starting from the three above-mentioned molecules, through recursive application of the graph transformation rules detailed in Appendix A.
Molecules with up to two carbon, four nitrogen, and four oxygen atoms were included in the generated network for comparability with [60]. The constructed reaction network had $44$ vertices and $116$ hyperedges, and shown in Figure 7. A list of molecules and reactions in the generated network is available in the GitHub data repository [data_repository]. To find probable pathways to glycine and glycolic acid, using small molecules MeCN (a hypothesized precursor to biotic molecules [61]), NH3, and H2O, an ILP was formulated to query the desired pathways. The ILP had $336$ variables and $1561$ constraints.
Figure 2: Top: A schematic depiction of the five best pathways to glyoxlic acid according to the chosen scoring function (energy levels not shown exactly to scale). The structures for the molecules in the corresponding vertices is listed in Table 3 and the weights assigned to the hyperedges listed in Table 4. Bottom: The energy profile of the best-scoring pathway (in blue) from above (energy levels shown to scale).
The five best scoring pathways to the target glycine is depicted in Figure 2. The best scoring pathway scores $143118$ on the chosen objective (III.2) and proceeds through iminoacetaldehyde ($v_{6}$), a pre-biotic intermediate hypothesized to be of significance [62, 63]. A water molecule acts as a catalyst in this pathway since it is used in reaction $e_{1}$ and generated back in reaction $e_{4}$. The second best pathway scores $143190$ on the objective (III.2). The last two reactions are the same for both (in fact, for all five) pathways. This second best pathway uses an ammonia molecule as a catalyst, molecules with a higher chemical potential (under unit concentrations) and reactions with higher energy barriers. The sum of the energy barriers for the reactions in this pathway is $200.02$ kJ mol-1 as opposed to $128.74$ kJ mol-1 for the best pathway. The third pathway branches off the second pathway and merges into the first, with a different reaction $e_{4}$. The fourth pathway uses a different higher energy molecule instead of iminoacetaldehyde. The fifth pathway has eight reactions, while the previous four had seven reactions. Despite having a lower sum of the energy barriers ($152.90$ kJ mol-1), it is ranked lower than the previous three pathways because it is longer. Figure 2 shows how interconnected the individual different pathways in the network might be, often with parallel branches and cross-connections, which leads to complex behavior of the system.
Figure 3: The six best pathways to glyoxlic acid according to the chosen scoring function. The structures for the molecules for the vertices is listed in Table 3 and the weights assigned to the hyperedges listed in Table 5.
Figure 3 shows the six best pathways to glycolic acid. The best scoring pathway has three reactions, the second best has four reaction, while the rest involve five reactions. Despite the third-listed pathway having a lower sum of energy barriers ($108.63$ kJ mol-1), it is ranked lower because it is longer, illustrating the effect of the second term in the objective (4).
V Discussions
In this submission, we have described a methodology for kinetically informed exploration of pathways in reaction networks. One part of the methodology is an automated workflow for annotating individual reactions with their energy barriers, computed using a diverse selection of tools— RDKit[64], xTB[48], ASE [56] and NeuralNEB [57]. The goal of this workflow is to minimize the laborious manual interventions and intensive expert knowledge often necessary when executing quantum chemical calculations to determine energy barriers. As demonstrated in Section IV, the workflow can be applied on networks of rather large scale, can work on the fly with generative network construction, and uses relatively moderate computational resources. The method does not take into account the interactions of solvent molecules while calculating the energy barriers. This and the question of how well models based on machine learning generalize past their training sets may lead to reasonable reservations on the precision of the estimates of the energy barriers for some use cases. However, our methodology offers the flexibility that this component can be substituted with alternatives, if desired.
As the other part of the proposed methodology, we formulated an ILP to query pathways in the annotated network and designed an objective function for pathway evaluation. The query formalism is at the same time both precise and very flexible in what queries can be formulated. The objective function is made under the simplifying assumption (borrowed from [31]) that all molecules in the reaction network are present with unit concentrations at all times. Additionally, the costs assigned to pathways depend on the underlying reaction network considered, so comparisons across networks are not directly possible. However, the question of how to evaluate pathways is complex [65, 66, 67] and the reader of course has the option to substitute a different objective function, based on other cost metrics. We do note, though, that our simplification and the formulation via ILP allows us to analyze much larger networks and using much less computational resources than alternatives like stochastic simulations or potential energy surface exploration. We also note that by a progressive constraining of the ILP, we in the end can return a list of high-scoring solutions, each representing structurally different pathways. Giving a list of good, structurally different solutions, rather than a single solution, helps mitigate disagreements about the details of pathway cost metrics.
Finally, we showcased the efficiency and utility of the described method on an actual reaction network, studying pathways to pre-biotically significant molecules, such as glycine and glycolic acid.
In summary, our work contributes to field of computationally assisted analysis and design of pathways in reaction networks represented as directed hypergraphs.
Acknowledgements.
This work was supported by generous funding from the European Union’s Horizon 2021 Research and Innovation program under Marie Sklodowska-Curie grant agreement no. 101072930 (TACsy — Training Alliance for Computational systems chemistry). The authors would also like to thank Dr. Maike Bergeler, Digitalization of R&D - Quantum Chemistry at BASF, Ludwigshafen am Rhein, Germany for discussions on the various methods to calculate the energy barriers for reactions.
Appendix A Graph Transformation rules used
The graph transformation rules used to expand the reaction network are listed in this appendix. A graph transformation rule consists of a subgraph on the left side which can be replaced by the subgraph on the right side when a match of the left-hand side is found in an already existing molecular graph in the network generated so far. Since these graph transformation rules represent a class of reactions, they can be considered as reaction templates.
[…]
Figure 4: Rule 1: Addition of water or ammonia to hydrogen cyanide (nitrile group) where $X\in\{\mathrm{N},\mathrm{O}\}$ to form an imine. The reverse of this rule was used while expanding the hypergraph.
[…]
Figure 5: Rule 2: Addition of water or ammonia to a double bond where $X,Y\in\{\mathrm{N},\mathrm{O}\}$ (in an imine or carbonyl group respectively). The reverse of this rule was also used while expanding the hypergraph.
[…]
Figure 6: Rule 3: 1,2 to 2,3 bond shift of the double bond to form tautomers where $X,Y\in\{\mathrm{N},\mathrm{O}\}$. The reverse of this rule was also used while expanding the hypergraph.
Appendix B List of Molecules in the presented pathways
In this appendix, we supply the molecular structure and the corresponding keys of the vertices (in the form $v_{i}$, where $i$ is some integer) in the pathways that are presented in Section IV. The two tables on the left side depict the molecules in the pathways presented in Figure 2, while the table on the right side depicts the molecules in the pathways presented in Figure 3.
Name Molecular Structure
$v_{0}$chemfigHO-[:30]-[:-30] [:-30]N
$v_{1}$chemfigNH_3
$v_{2}$chemfigH_2O
$v_{3}$chemfigHO-[:30]-[:-30](-[:-90]OH)=[:30]NH
$v_{4}$chemfigHO-[:30]=[:-30](-[:-90]OH)-[:30]NH_2
$v_{5}$chemfigO=[:30]-[:-30](-[:-90]OH)-[:30]NH_2
$v_{6}$chemfigO=[:30]-[:-30]=[:30]NH
$v_{7}$chemfigHO-[:30](-[:90]OH)-[:-30]=[:30]NH
$v_{8}$chemfigHO-[:30](-[:90]OH)=[:-30]-[:30]NH_2
$v_{9}$chemfigHO-[:30](=[:90]O)-[:-30]-[:30]NH_2
Name Molecular Structure
$v_{13}$chemfigHO-[:30]-[:-30](-[:-90]NH_2)=[:30]NH
$v_{14}$chemfigHO-[:30]=[:-30](-[:-90]NH_2)-[:30]NH_2
$v_{15}$chemfigO=[:30]-[:-30](-[:-90]NH_2)-[:30]NH_2
$v_{16}$chemfigHO-[:30](-[:90]OH)-[:-30](-[:-90]NH_2)-[:30]NH_2
$v_{36}$chemfigH_2N-[:30](-[:90]OH)-[:-30](-[:-90]OH)-[:30]NH_2
$v_{52}$chemfigH_2N-[:30](=[:90]O)-[:-30]-[:30]OH
Name Molecular Structure
$v_{0}$chemfigHO-[:30]-[:-30] [:-30]N
$v_{1}$chemfigNH_3
$v_{2}$chemfigH_2O
$v_{3}$chemfigHO-[:30]-[:-30](-[:-90]OH)=[:30]NH
$v_{6}$chemfigHO-[:30]-[:-30](-[:-60]OH)(-[:-120]HO)-[:30]NH_2
$v_{7}$chemfigHO-[:30]-[:-30](=[:-90]O)-[:30]OH
$v_{15}$chemfigHO-[:30]-[:-30](=[:-90]O)-[:30]NH_2
$v_{24}$chemfigHO-[:30]=[:-30](-[:-90]OH)-[:30]NH_2
$v_{33}$chemfigHO-[:30]-[:-30](-[:-90]NH_2)=[:30]NH
$v_{34}$chemfigHO-[:30]-[:-30](-[:-60]NH_2)(-[:-120]HO)-[:30]NH_2
Table 3: The molecular structures corresponding to the the labeled vertices in the pathways in Figure 2 (two left columns) and the pathways in Figure 3 (right column).
Appendix C Energy barriers in the presented pathways
The estimated energy barriers for the reactions in the generated reaction network, which are assigned as weights to the hyperedges during the pathway search and scoring process, are presented in the two tables in this appendix. Table 4 consists of five columns that list the assigned weights for the hyperedges in the pathways shown in Figure 2, while Table 5 contains six columns that list the assigned weights for the hyperedges in the pathways illustrated in Figure 3.
The shaded cells indicate hyperedges that have already been used in a previously enumerated pathway. For instance, all five pathways in Figure 2 share the final two reactions. Consequently, the next four pathways reuse the last two reactions of the first pathway, and the last two rows for the corresponding columns are shaded to reflect this overlap.
Reaction Energy barriers (in kJ mol-1)
count          pathway 1          pathway 2          pathway 3          pathway 4          pathway 5
$e_{1}$ $13.95$ $33.18$ $33.18$ $13.95$ $13.95$
$(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{1},v_{13})$ $(v_{0}+v_{1},v_{13})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{2},v_{3})$
$e_{2}$ $7.67$ $64.17$ $64.17$ $7.67$ $18.67,13.15$
$(v_{3},v_{4})$ $(v_{13},v_{14})$ $(v_{13},v_{14})$ $(v_{3},v_{4})$ $(v_{3},v_{52},v_{4})$
$e_{3}$ $57.69$ $28.40$ $28.40$ $57.69$ $57.69$
$(v_{4},v_{5})$ $(v_{14},v_{15})$ $(v_{14},v_{15})$ $(v_{4},v_{5})$ $(v_{4},v_{5})$
$e_{4}$ $1.67$ $19.86$ $29.35$ $67.36$ $1.67$
$(v_{5},v_{6}+v_{2})$ $(v_{15}+v_{2},v_{16})$ $(v_{15},v_{6}+v_{1})$ $(v_{5}+v_{2},v_{36})$ $(v_{5},v_{6}+v_{2})$
$e_{5}$ $17.14$ $23.79$ $17.14$ $60.84$ $17.14$
$(v_{6}+v_{2},v_{7})$ $(v_{16},v_{7}+v_{1})$ $(v_{6}+v_{2},v_{7})$ $(v_{36},v_{7}+v_{2})$ $(v_{6}+v_{2},v_{7})$
$e_{6}$ $0.19$ $0.19$ $0.19$ $0.19$ $0.19$
$(v_{7},v_{8})$ $(v_{7},v_{8})$ $(v_{7},v_{8})$ $(v_{7},v_{8})$ $(v_{7},v_{8})$
$e_{7}$ $30.42$ $30.42$ $30.42$ $30.42$ $30.42$
$(v_{8},v_{9})$ $(v_{8},v_{9})$ $(v_{8},v_{9})$ $(v_{8},v_{9})$ $(v_{8},v_{9})$
sum $128.74$ $200.02$ $202.85$ $238.13$ $152.90$
Table 4: From the left, the columns list the weights assigned to the hyperedges along with the sum of the weights, for the five pathways in Figure 2.
Reaction Energy barrier
count          pathway 1          pathway 2          pathway 3          pathway 4          pathway 5          pathway 6
$e_{1}$ $13.95$ $13.95$ $13.95$ $33.18$ $13.95$ $33.18$
$(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{1},v_{33})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{1},v_{3}3)$
$e_{2}$ $7.67$ $48.612$ $80.53$ $48.61$
$(v_{3},v_{24})$ $(v_{33}+v_{2},v_{34})$ $(v_{3}+v_{1},v_{34})$ $(v_{33}+v_{2},v_{34})$
$e_{3}$ $26.61$ $18.68$ $11.76$ $43.64$ $43.64$ $33.08$
$(v_{3}+v_{2},v_{6})$ $(v_{3},v_{15})$ $(v_{24},v_{15})$ $(v_{34},v_{15}+v_{1})$ $(v_{34},v_{15}+v_{1})$ $(v_{34},v_{3}+v_{1})$
$e_{4}$ $0.62$ $0.62$ $0.62$ $0.62$ $0.62$
$(v_{15}+v_{2},v_{6})$ $(v_{15}+v_{2},v_{6})$ $(v_{15}+v_{2},v_{6})$ $(v_{15}+v_{2},v_{6})$ $(v_{3}+v_{2},v_{6})$
$e_{5}$ $74.62$ $74.62$ $74.62$ $74.62$ $74.62$ $74.62$
$(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$
sum $118.19$ $107.87$ $108.63$ $200.68$ $213.38$ $219.11$
Table 5: From the left, the columns list the weights assigned to the hyperedges along with the sum of the weights, for the six pathways in Figure 3.
Appendix D Constructed Reaction Network
Figure 7: Hypergraph with $|V|=44$ and $|E|=116$, representing the reaction network considered in this work.